# Local converse theorems and Langlands parameters

### v5

@@ 2. Preliminaries
All representations that we consider have complex coefficients. The main reference for the material in this section are [Tat] and [Bor], and [DHKM, Section 6]. Let $F$ be a non Archimedean local field with residual cardinality $q$, Weil group $\mathrm{W}_{F}$ and Weil-Deligne group ${\mathrm{WD}}_{F}=\mathrm{W}_{F}\times{\mathrm{SL}}_{2}({\mathbb{C}})$. We do not impose any restriction on the characteristic of $F$. We denote by $I_{F}$ the inertia subgroup of $\mathrm{W}_{F}$, and fix ${\mathrm{Fr}}_{F}$ a geometric Frobenius element of $\mathrm{W}_{F}$. We set $\nu_{F}:\mathrm{W}_{F}\to{\mathbb{C}}^{\times}$ to be the unique character trivial on $I_{F}$ such that $\nu_{F}({\mathrm{Fr}}_{F})=q^{-1}$. Via local class field theory, the map $\nu$ corresponds to the normalized absolute value |||\ || | on $F^{\times}$. For $\gamma\in\mathrm{W}_{F}$ we denote by $d(\gamma)$ the integer such that $\nu(\gamma)=q^{-d(\gamma)}$. Let $G$ be the group of $F$-points of a quasi-split connected reductive group defined over a local field $F$. We denote by $\widehat{G}$ its Langlands dual group and its $L$-group by ${}^{L}G:=\widehat{G}\rtimes\mathrm{W}_{F}$, where $\mathrm{W}_{F}$ acts trivially on $\widehat{G}$ if $G$ is split over $F$. For example ${}^{L}{\mathrm{GL}}_{n}(F)={\mathrm{GL}}_{n}({\mathbb{C}})\times\mathrm{W}_{F}$. By definition, what we call a representation of ${}^{L}G$ is a finite dimensional semi-simple and continuous representation.
@@ 2.1. Langlands parameters and their local constants
We will also need another version of admissible homomorphisms, which is that of Weil-Deligne parameters as in [DHKM, Definition 6.1]. We denote by $\widehat{\mathfrak{g}}$ the Lie algebra of $\widehat{G}$, and by $\mathrm{Ad}:{}^{L}G\to{\mathrm{GL}}(\widehat{\mathfrak{g}})$ the adjoint representation of ${}^{L}G$. A Weil-Deligne parameter consists of a pair $(\tau,N)$, where $\tau:\mathrm{W}_{F}\to{}^{L}G$ is semi-simple homomorphism, and $N$ is a nilpotent element in $\widehat{\mathfrak{g}}$ satisfying
@@ 2.2. Generic parameters
[…]
@@ 4. Acceptable reductive algebraic groups over ${\mathbb{C}}$
On the other hand even orthogonal ${\mathrm{SO}}_{2n}({\mathbb{C}})$ for $n$ larger than 333, are for example unacceptable.
@@ 5. A consequence of results of Silberger and Zink
In this section we consider $\widehat{G}$ as in Section 4. Our goal is to prove Theorem 5.3, which states that acceptability of $\widehat{G}$ implies its ”$\mathrm{W}_{F}$-acceptability”. We refer to [SZ, Section 5.1] for the notions of hyperbolic and elliptic elements in a complex reductive group ([SZ] more generally considers L-groups of quasi-split groups which might be disconnected, but we will not need this degree of generality here). If $\varphi:\mathrm{W}_{F}\to\widehat{G}$ is a semi-simple homomorphism (identified with an admissible homomorphism of ${\mathrm{WD}}_{F}$ trivial on ${\mathrm{SL}}_{2}({\mathbb{C}})$) and $\gamma\in\mathrm{W}_{F}$, we also refer to [SZ, Section 5.1] for the definition of the hyperbolic and elliptic parts of $\varphi(\gamma)$. Now we state a consequence of [SZ, Proposition 5.3].
@@ 6. A local converse theorem for Langlands parameters
[…]
@@
[…]
@@ 7. Unacceptable parameters
Suppose that $4|q-1$ and that $n\geq 3$. Then there exist non equivalent generic homomorphisms $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}{\mathrm{SO}}_{2n}(F)$ such that for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
@@
Under the condition that $4|q-1$, and by class field theory, the group $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ is a quotient of $\mathrm{W}_{F}$ because it is a quotient of $F^{\times}$. In particular [Wei, Theorem 24] provides locally but non globally conjugate parameters $\varphi_{i}:\mathrm{W}_{F}\to{}^{L}{\mathrm{SO}}_{2n}(F)$, $i=1,2$, by inflation. Let us slightly rephrase and explicate Weidner’s example. We denote by $w_{n}$ the antidiagonal matrix in ${\mathrm{GL}}_{n}({\mathbb{C}})$ with ones on the second diagonal. We realize ${}^{L}{\mathrm{SO}}_{2n}(F)={\mathrm{SO}}_{2n}({\mathbb{C}})$ as the group
@@
where $\varpi_{F}$ is a uniformizer of $F$, $O_{F}$ is the ring of integers of $F$, and $u$ has order $q-1$. Then
@@
[…]
@@
This is impossible since $\mathrm{diag}(i,-i)$ is not ${\mathrm{SO}}_{2}({\mathbb{C}})$-conjugate to $\mathrm{diag}(-i,i)$, as ${\mathrm{SO}}_{2}({\mathbb{C}})$ consists of diagonal matrices with determinant one. Moreover the centralizers of $\varphi_{1}$ and $\varphi_{2}$ are both equal to the set
$\{\mathrm{diag}(z,A,B,w_{n-2}A^{-T}w_{n-2},z^{-1}),\ A\in{\mathrm{GL}}_{n-2}({\mathbb{C}}),\ B\in{\mathrm{SO}}_{2}({\mathbb{C}}),\ z\in{\mathbb{C}}^{\times}\}.$
In particular, according to [DHKM, Proposition 6.10], we can take a nilpotent element $N$ in the Lie algebra of ${\mathrm{SO}}_{2n}({\mathbb{C}})$ such that $(\varphi_{1},N)$ and $(\varphi_{2},N)$ are Weil-Deligne parameters, which are moreover generic. We then take $\phi_{i}$ corresponding to $(\varphi_{i},N)$, and apply Proposition 7.2. ∎
@@
The monodromy operator $N$ in the proof above must be equal to zero since the eigenvalues of the adjoint action of $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ on the Lie algebra of $\widehat{G}$ are of absolute value one, whereas the image of the Frobenius element in the group $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ would act with eigenvalue $q^{-1}$ on $N$ if $N$ was nonzero.
@@ 8. A variant
The statement is obviously true in the first subcase of Case (a), thanks to Theorem 3.3 and Lemma 2.1. For unitary groups we recall that $\varphi_{1}\sim\varphi_{2}$ if and only if ${\varphi_{1}}_{|\mathrm{W}_{E}}\sim{\varphi_{2}}_{|\mathrm{W}_{E}}$ ([Pra, Proposition 7.10]). On the other hand, under the assumptions of Statement 8.1, we deduce from Theorem 3.3 that ${\varphi_{1}}_{|\mathrm{W}_{E}}\sim{\varphi_{2}}_{|\mathrm{W}_{E}}$, hence the result now follows from Lemma 2.1. In Case (b), the normalizer of $r(\mathrm{G}_{2}({\mathbb{C}}))$ is equal to ×${\mathbb{C}}^{\times}\mathrm{G}_{2}({\mathbb{C}})$. This follows for example from applying [Gri, Theorem 1.1], and observing that the centralizer of $r(\mathrm{G}_{2}({\mathbb{C}}))$ is the center of ${\mathrm{GL}}_{7}({\mathbb{C}})$ according to Schur’s lemma. Then the theorem in Case (b) follows from Lemma 2.1, Theorem 3.3 and [Gri, Theorem 1.1]. In cases (c), (d), (e) and (f), the normalizers are respectively ×${\mathbb{C}}^{\times}{\mathrm{SO}}_{2n+1}({\mathbb{C}})$, ×${\mathbb{C}}^{\times}{\mathrm{Sp}}_{2n}({\mathbb{C}})$ and ×${\mathbb{C}}^{\times}{\mathrm{O}}_{2n+2}({\mathbb{C}})$, as follows for example from [Gri, Section 2]. The result that we claim follows from Lemma 2.1, Theorem 3.3 and [Gri, Theorem 2.3]. ∎
@@
It is false in general that one can replace the normalizer of $r({}^{L}G)$ by $r({}^{L}G)$ itself. For example in Case (f) above, two admissible homomorphisms $\phi_{1}$ and $\phi_{2}$ with $\phi_{2}=s\phi_{1}s^{-1}$ for orthogonal $s$ with determinant 1-1- 1 will satisfy the equality of twisted gamma factors, but might be inequivalent.

### v6

@@ 2. Preliminaries
All representations that we consider have complex coefficients. The main reference for the material in this section are [Tat] and [Bor], and [DHKM, Section 6]. Let $F$ be a non Archimedean local field with residual cardinality $q$, Weil group $\mathrm{W}_{F}$ and Weil-Deligne group ${\mathrm{WD}}_{F}=\mathrm{W}_{F}\times{\mathrm{SL}}_{2}({\mathbb{C}})$. We do not impose any restriction on the characteristic of $F$. We denote by $I_{F}$ the inertia subgroup of $\mathrm{W}_{F}$, and fix ${\mathrm{Fr}}_{F}$ a geometric Frobenius element of $\mathrm{W}_{F}$. We set $\nu_{F}:\mathrm{W}_{F}\to{\mathbb{C}}^{\times}$ to be the unique character trivial on $I_{F}$ such that $\nu_{F}({\mathrm{Fr}}_{F})=q^{-1}$. Via local class field theory, the map $\nu$ corresponds to the normalized absolute value |||\ | on $F^{\times}$. For $\gamma\in\mathrm{W}_{F}$ we denote by $d(\gamma)$ the integer such that $\nu(\gamma)=q^{-d(\gamma)}$. Let $G$ be the group of $F$-points of a quasi-split connected reductive group defined over a local field $F$. We denote by $\widehat{G}$ its Langlands dual group and its $L$-group by $^{L}G:=\widehat{G}\rtimes\mathrm{W}_{F}$, where $\mathrm{W}_{F}$ acts trivially on $\widehat{G}$ if $G$ is split over $F$. For example $^{L}{\mathrm{GL}}_{n}(F)={\mathrm{GL}}_{n}({\mathbb{C}})\times\mathrm{W}_{F}$. By definition, what we call a representation of $^{L}G$ is a finite dimensional semi-simple and continuous representation.
@@ 2.1. Langlands parameters and their local constants
We say that a continuous homomorphism $\varphi:\mathrm{W}_{F}\to{}^{L}G$ is semi-simple if the projection of its image on $\widehat{G}$ consists of semi-simple elements, and the second coordinate of $\phi$ is the identity of $\mathrm{W}_{F}$. We say that a homomorphism $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ is admissible if $\phi_{|\mathrm{W}_{F}}$ is semi-simple, and $\phi_{|{\mathrm{SL}}_{2}({\mathbb{C}})}$ is algebraic. Let $X$ be a group and $\tau_{1},\ \tau_{2}:X\to{}^{L}G$ two homomorphisms, we write $\tau_{1}\underset{L}{\sim}\tau_{2}$ if they are conjugate under $^{L}G$. We say that $\tau_{1}$ and $\tau_{2}$ are equivalent and write $\tau_{1}\sim\tau_{2}$ if they are conjugate under $\widehat{G}$. An admissible homomorphism $\phi:{\mathrm{WD}}_{F}\to{}^{L}{\mathrm{GL}}_{n}(F)$ identifies with a (smooth) semi-simple representation of ${\mathrm{WD}}_{F}$ of dimension $n$, because composition with the projection onto ${\mathrm{GL}}_{n}({\mathbb{C}})$ is injective, and we will do this identification freely.
@@ 2.2. Generic parameters
Let $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}G$ be two generic homomorphisms, then $\phi_{1}\sim\phi_{2}\iff\varphi_{1}\sim\varphi_{2}$. Similarly $\phi_{1}\underset{L}{\sim}\phi_{2}\iff\varphi_{1}\underset{L}{\sim}\varphi_{2}$.
@@ 4. Acceptable reductive algebraic groups over $\mathbb{C}$}
On the other hand even orthogonal ${\mathrm{SO}}_{2n}({\mathbb{C}})$ for $n$ larger than $3$, are for example unacceptable.
@@ 5. A consequence of results of Silberger and Zink
In this section we consider $\widehat{G}$ as in Section 4. Our goal is to prove Theorem 5.3, which states that acceptability of $\widehat{G}$ implies its $\mathrm{W}_{F}$-acceptability”. We refer to [SZ, Section 5.1] for the notions of hyperbolic and elliptic elements in a complex reductive group ([SZ] more generally considers L-groups of quasi-split groups which might be disconnected, but we will not need this degree of generality here). If $\varphi:\mathrm{W}_{F}\to\widehat{G}$ is a semi-simple homomorphism (identified with an admissible homomorphism of ${\mathrm{WD}}_{F}$ trivial on ${\mathrm{SL}}_{2}({\mathbb{C}})$) and $\gamma\in\mathrm{W}_{F}$, we also refer to [SZ, Section 5.1] for the definition of the hyperbolic and elliptic parts of $\varphi(\gamma)$. Now we state a consequence of [SZ, Proposition 5.3].
@@ 6. A local converse theorem for Langlands parameters
    Then $\varphi_{1}\underset{L}{\sim}\varphi_{2}$. In particular $\phi_{1}\underset{L}{\sim}\phi_{2}$ if both $\phi_{1}$ and $\phi_{2}$ are generic (see Lemma 2.1). If $G$ is $F$-split one can replace $\underset{L}{\sim}$ by $\sim$ in the two previous sentences.
@@
We set $J=\begin{pmatrix}&&&1\\ &&-1&\\ &\iddots&&\\ (-1)^{n-1}\end{pmatrix}$. We have $\widehat{G}={\mathrm{GL}}_{n}({\mathbb{C}})$ and for $g\in{\mathrm{GL}}_{n}({\mathbb{C}})$ we write $g^{-T}$ for the inverse of the transpose of $g$. The L-group $^{L}G$ admits ${\mathrm{GL}}_{n}({\mathbb{C}})\rtimes{\mathbb{Z}}/2{\mathbb{Z}}$ as a quotient, where the non trivial element $\alpha$ of ${\mathbb{Z}}/2{\mathbb{Z}}$ by the automorphism $\alpha(g)=Jg^{-T}J^{-1}$ of ${\mathrm{GL}}_{n}({\mathbb{C}})$. Let $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}G$ be stable and discrete homomorphisms. We will prove that if $r\circ\phi_{1}$ and $r\circ\phi_{2}$ are equivalent representations of $\mathrm{W}_{F}$ for all representations $r$ of $^{L}G$, then $\phi_{1}\underset{L}{\sim}\phi_{2}$. We take $\mathrm{St}$ the identity representation of ${\mathrm{GL}}_{n}({\mathbb{C}})$, and set
@@ 7. Unacceptable parameters
Suppose that 4|q14|q-1 and that $n\geq 3$. Then there exist non equivalent generic homomorphisms $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}{\mathrm{SO}}_{2n}(F)$ such that for any irreducible representation $r$ of $^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
@@
Under the condition that 4|q14|q-1, and by class field theory, the group $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ is a quotient of $\mathrm{W}_{F}$ because it is a quotient of $F^{\times}$. In particular [Wei, Theorem 24] provides locally but non globally conjugate parameters $\varphi_{i}:\mathrm{W}_{F}\to{}^{L}{\mathrm{SO}}_{2n}(F)$, $i=1,2$, by inflation. Let us slightly rephrase and explicate Weidner’s example. We denote by $w_{n}$ the antidiagonal matrix in ${\mathrm{GL}}_{n}({\mathbb{C}})$ with ones on the second diagonal. We set $G={\mathrm{SO}}_{2n}(F)$ and realize $\widehat{G}={\mathrm{SO}}_{2n}({\mathbb{C}})$ as the group
@@
where $\varpi_{F}$ is a uniformizer of $F$, $O_{F}$ is the ring of integers of $F$, and $u$ has order q1q-1. Then
@@
  • when $b$ is odd and $a$ is even then $\lambda_{1}(\varpi_{F}^{a}u^{b})$ and $\lambda_{2}(\varpi_{F}^{a}u^{b})$ are conjugate by the permutation matrix corresponding to the double transposition $(n\ n+1)\circ(1\ 2n)$,
@@
This is impossible since $\mathrm{diag}(i,-i)$ is not ${\mathrm{SO}}_{2}({\mathbb{C}})$-conjugate to $\mathrm{diag}(-i,i)$, as ${\mathrm{SO}}_{2}({\mathbb{C}})$ consists of diagonal matrices with determinant one. Finally, by [DHKM, Proposition 6.10], there exists a monodromy operator $N_{i}\in\mathcal{N}_{\varphi_{i}}$ making $(\varphi_{i},N_{i})$ a generic Weil-Deligne parameter. We then take $\phi_{i}$ corresponding to $(\varphi_{i},N_{i})$, and apply Proposition 7.2. ∎
@@
The monodromy operators $N_{i}$ in the proof above must be equal to zero since the eigenvalues of the adjoint action of $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ on the Lie algebra of $\widehat{G}$ are of absolute value one, whereas the image of the Frobenius element in the group $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ would act with eigenvalue $q^{-1}$ on $N$ if $N$ was nonzero.
@@ 8. A variant
The statement is obviously true in the first subcase of Case (a), thanks to Theorem 3.3 and Lemma 2.1. For unitary groups we recall that $\varphi_{1}\sim\varphi_{2}$ if and only if ${\varphi_{1}}_{|\mathrm{W}_{E}}\sim{\varphi_{2}}_{|\mathrm{W}_{E}}$ ([Pra, Proposition 7.10]). On the other hand, under the assumptions of Statement 8.1, we deduce from Theorem 3.3 that ${\varphi_{1}}_{|\mathrm{W}_{E}}\sim{\varphi_{2}}_{|\mathrm{W}_{E}}$, hence the result now follows from Lemma 2.1. In Case (b), the normalizer of $r(\mathrm{G}_{2}({\mathbb{C}}))$ is equal to ${\mathbb{C}}^{\times}\mathrm{G}_{2}({\mathbb{C}})$. This follows for example from applying [Gri, Theorem 1.1], and observing that the centralizer of $r(\mathrm{G}_{2}({\mathbb{C}}))$ is the center of ${\mathrm{GL}}_{7}({\mathbb{C}})$ according to Schur’s lemma. Then the theorem in Case (b) follows from Lemma 2.1, Theorem 3.3 and [Gri, Theorem 1.1]. In cases (c), (d), (e) and (f), the normalizers are respectively ${\mathbb{C}}^{\times}{\mathrm{SO}}_{2n+1}({\mathbb{C}})$, ${\mathbb{C}}^{\times}{\mathrm{Sp}}_{2n}({\mathbb{C}})$ and ${\mathbb{C}}^{\times}{\mathrm{O}}_{2n+2}({\mathbb{C}})$, as follows for example from [Gri, Section 2]. The result that we claim follows from Lemma 2.1, Theorem 3.3 and [Gri, Theorem 2.3]. ∎
@@
It is false in general that one can replace the normalizer of $r({}^{L}G)$ by $r({}^{L}G)$ itself. For example in Case (f) above, two admissible homomorphisms $\phi_{1}$ and $\phi_{2}$ with $\phi_{2}=s\phi_{1}s^{-1}$ for orthogonal $s$ with determinant 1-1 will satisfy the equality of twisted gamma factors, but might be inequivalent.